### v1

# Group Testing with General Correlation Using Hypergraphs

Abstract
Group testing, a problem with diverse applications across multiple disciplines, traditionally assumes independence across nodes’ states. Recent research, however, focuses on real-world scenarios that often involve correlations among nodes, challenging the simplifying assumptions made in existing models. In this work, we consider a comprehensive model for arbitrary statistical correlation among nodes’ states. To capture and leverage these correlations effectively, we model the problem by hypergraphs, inspired by [GLS22], augmented by a probability mass function on the hyper-edges.
Using this model, we first design a novel greedy adaptive algorithm capable of conducting informative tests and dynamically updating the distribution. Performance analysis provides upper bounds on the number of tests required, which depend solely on the entropy of the underlying probability distribution and the average number of infections. We demonstrate that the algorithm recovers or improves upon all previously known results for group testing settings with correlation. Additionally, we provide families of graphs where the algorithm is order-wise optimal and give examples where the algorithm or its analysis is not tight.
We then generalize the proposed framework of group testing with general correlation in two directions, namely noisy group testing and semi-non-adaptive group testing. In both settings, we provide novel theoretical bounds on the number of tests required.
1 Introduction
Group testing is a classical problem that has been extensively studied in the fields of computer science [DHH00, DBGV05, SG59, PR08] and information theory [AJS19, ABJ14, Wol85, CJSA14], focusing on efficiently identifying a small number of defective items within a large population by grouping items together in tests. Originally motivated by the need to screen for syphilis during World War II, group testing has recently garnered renewed attention due to its applications in areas such as COVID-19 testing [GG20, NHL20], privacy-preserving [ICDÖ23], and DNA sequencing [Dan23].
Group testing involves a population (set of nodes) where a subset is infected, and any sub-population can be tested. A positive test result indicates at least one infection within the subset, and the primary goal is to identify all infected individuals. Historically, group testing has been studied under the assumption of independent node states (see [LCH+14, DHH00] and the references therein). However, recent advancements have highlighted the oversimplification of this assumption. For example, in disease spread scenarios, individuals within the same household exhibit correlation in their risk; if one member is infected, the likelihood of infection among other members increases [NSG+21a, NKC+22, AU23]. To a lesser extent, a weaker correlation exists in the population of a neighbor where an outbreak might occur, and even weaker correlation exists in the population of a city, so different strengths of correlation can be assumed in a general population. More generally, the physical and biological interactions that govern disease propagation impose a correlation structure on the states of the nodes. Similarly, in network fault diagnosis, geographically localized connections/devices (to which we refer to as nodes) are more likely to experience simultaneous faults due to their reliance on shared physical infrastructure and exposure to common external factors, such as environmental conditions or power supply issues [XAAA23]. Furthermore, the cascading nature of many failure mechanisms introduces correlations across the states of these nodes, as a fault in one part of the system can increase the likelihood of faults in nearby components. Examples include an outage of power in some part of the network[BBH+14], an overload in a specific district[AW17], or typically localized cascading line failures in the transmission system of the power grid [SMZ14]. Modeling these correlations and exploiting them in the design of group testing algorithms can enhance accuracy and efficiency, leading to more effective group testing.
To capture correlation, much of the existing literature propose models that are context-specific and oftentimes oversimplified. Furthermore, the proposed group testing methods are tied to the underlying correlation models that are studied. In the context of disease propagation, the correlation models are built either on the notion of proximity (defined through graphs) [NKC+22, AU23, LSS22, AU21] or in conjunction with a disease spread model [ZH23, BMR21, NSG+23, ACÖ23]. In the context of network fault diagnosis, the correlation is based on physical proximity between the nodes in a network, or frequency of communication between two nodes [XAAA23].
In this work, we propose a unifying framework for group testing in the presence of correlation, without being restricted by context-specific spread models. Our objective is to study a general correlation model, i.e., a general joint distribution denoted by $\mathcal{D}$ on the state of the nodes, where $\mathcal{D}(S)$ represents the probability that exactly a subset $S$ of individuals are infected. Our aim is to devise group testing algorithms tailored to this expansive scope without substantial simplification. Our proposed framework, along with the group testing strategies we propose, can directly apply to any application with an underlying statistical correlation model.
The paper is organized as follows. In the remainder of this section, we discuss related work and the contributions of this paper. In Section 2, we introduce our proposed model and the mathematical problem formulation. In Section 3, we study adaptive group testing and propose a greedy adaptive algorithm that imposes no limitation on the underlying distribution of the hyper-edges. The algorithm is analyzed in Section 4. In Section 5, we apply our algorithm to prior models to recover and improve prior results. Section 6 provides examples where the algorithm is order-optimal and others where it is not. In Section 7, we present algorithms that are not adaptive, and in Section 8, we consider the setting where tests are noisy. Finally, we conclude and discuss open problems in Section 9.
1.1 Related Work
Traditionally, group testing has been examined via two primary paradigms; namely, combinatorial group testing and probabilistic group testing. Here, we first go through the combinatorial version and then introduce the probabilistic version. In each, we delve into recent research efforts that establish correlation models and harness this correlation to reduce the number of required tests.
Combinatorial group testing considers scenarios in which, out of a population of size $n$, a maximum of $d$ individuals are infected [DHH00]. Adaptive group testing, where each test is designed based on the result of the previous tests, has been proved to efficiently identify the infected set using at most $d\log n+O(d)$ tests, complemented by a corresponding lower bound. In the non-adaptive case, where the tests are designed ahead of time, the best known upper bound is $d^{2}\log(n/d)$, complemented with an almost matching lower bound of $\Omega(\frac{d^{2}\log n}{\log d})$ [EFF85]. There has also been recent literature on group testing with only a few rounds of adaptation [WGG24, COGHKL20, Sca19]. For example, [WGG24] shows that using $O(\log d)$ rounds of testing, almost all infected nodes can be recovered using almost $\frac{1}{\text{capacity}(Z)}d\log n/d$ tests where the tests are noisy and $Z$ is a generic channel corrupting the test outcomes. More recent works have aimed not only to be order-wise optimal but also to determine the exact constant coefficients for the number of tests required [CJSA14, SC16, AJS16, WGG24].
Another line of work is quantitative group testing, which it explores scenarios where the test results are not binary but instead, a test returns the exact number of infections within a group. It is shown that in the adaptive setting, $O(\log n)$ tests are enough [Chr80, Aig86] and later the coefficient of $\log n$ was improved[GMSV92]. Recently, [SJ24b] designed low-complexity algorithms with efficient construction and decoding. For non-adaptive group testing, [HKM22] obtains $O(d\log n)$ which matches the information-theoretic lower bounds up to a factor. [SJ24a] has improved the decoding complexity.
In a recent study, [GLS22] studies a scenario in which only specific subsets of individuals can be infected, collectively referred to as the set of candidate subsets denoted as $E$. This model can partly capture the correlation between the nodes. [GLS22] models candidate subsets as hyperedges of a hypergraph defined on the set of individuals (nodes) $V$ and assumes that each hyperedge has a maximum size of $d$. It provides non-adaptive and adaptive group testing strategies that require $O(d\log E)$ and $O(\log|E|+d^{2})$ tests, respectively. These bounds are complemented with the lower bound $\log|E|+d$ for a specific “worst-case” hypergraph. This work is closest to our work, both in terms of the underlying model and the conceptual framework for algorithm design. However, their framework does not allow different likelihoods of the edges. This assumption is not practical for many applications. For instance, very large and very small edges might be unlikelier than the medium-sized edges. Or different classes of individuals may have different risk factors. Notably, our approach encompasses not only the definition of hyperedges but also introduces a distribution over them to introduce likelihood, and as such our work stands in the so called probabilistic group testing paradigm.
There is a variant of group testing, also known as (hyper)graph learning, which is based on graphs and hypergraphs where items are semi-defective, meaning that the test result is positive when specific subsets of the (hyper)graph are included in the test [Aba18, AN19, BS24]. Several bounds have been established for general graphs and hypergraphs [Aba18, AN19], and these bounds have been improved by assuming specific structures on the hypergraphs [BS24]. Although these works utilize the concepts of graphs and hypergraphs, our problem differs in its fundamental nature.
There have been other works where specific structures of graphs and hypergraphs are assumed and the testing is designed by the structure to improve the testing even further. For instance, in [BCN21] and [BS23] they assume that the graph is a line and a block of consecutive nodes is positive. Another example is [NKC+22] where they give theoretical bounds on the number of tests when the graph is a tree or grid. In this work, we do not have any constraints on the structure of the hypergraph.
Probabilistic group testing was studied in [LCH+14], where each individual $i$ is infected with a probability $p_{i}$, independently of others. Their research led to the establishment of near-optimal bounds for both adaptive and non-adaptive group testing, quantified as $O(H(X)+\mu)$, where $\mu=\sum_{i}p_{i}$ represents the average number of infections, and $H(X)=\sum_{i}-p_{i}\log p_{i}$ denotes the entropy. In recent years, there has been a surge of interest in modeling correlations in group testing scenarios. Several noteworthy developments in this area include: modeling correlation using edge-faulty graphs [NKC+22, AU23] where a simple graph $G$ is considered as the underlying contact graph and each edge is dropped with some probability, forming random components. The nodes within a component are then assumed to be in the same state. In [NKC+22], it was shown for several families of graphs that exploiting correlation can lead to significant improvement in the required number of tests (compared to independent group testing). [AU23] shows improvement when the realized graphs have a certain structure, i.e. they are nested. Another line of work is using communities to model correlation [NSG+21a, ACO21, NSG+23, NSG+21b, JCM24]. For example, in [NSG+21a], it is assumed that each community is infected with some probability, and if a community is infected, all of its members are infected with some non-zero probability independent of the others. If a community is not infected, none of its members are infected. For this model, the authors designed efficient adaptive and non-adaptive algorithms that significantly reduce the number of tests by accounting for the structure and pooling the infected communities. They also developed more reliable tests using the structure when the tests are noisy. [ACO21], considers similar communities, but the communities become infected by random initial seeds. The authors design an adaptive algorithm that significantly reduces the number of tests, further supporting the benefits of exploiting correlation.
1.2 Contributions
The contributions of this work are as follows:
  • We model arbitrary statistical correlation by employing hypergraphs, drawing inspiration from [GLS22]. A hypergraph $G=(V,E)$ has node set $V$, $|V|=n$, and edge set $E$, where each hyperedge $e\in E$ is a subset of nodes: $e\subseteq V$. Enhancing the combinatorial model in [GLS22], we consider an arbitrary probability distribution of infection defined over the hyperedges. This captures arbitrary statistical correlation among infection of different nodes.
  • We propose an adaptive group testing algorithm that is capable of exploiting correlation by the means of updating the posterior distribution on the hyperedges given the previous test results. The algorithm works in two stages. Stage 1 focuses on conducting “informative” tests, which significantly narrows down the search space. When these tests are no longer feasible, the algorithm transitions to a second stage where it tests the remaining uncertain nodes individually. We demonstrate that this algorithm successfully identifies the infected set and requires an expected number of tests that can be upper bounded as a function of entropy $H(X)$ and the expected number of infections $\mu$. For distributions that are concentrated around $\mu$, the bound takes the form of ${O}(H(X)+\mu)$.
  • We provide various examples to illustrate the extent of optimality of the algorithm. $H(X)$ is a known lower bound on the number of tests. We show that our proposed algorithm, and the upper bound we establish on the number of tests, ${O}(H(X)+\mu)$, is worst-case order-optimal. We further show that (i) $H(X)$ can be a loose lower bound with a significant gap to the number of tests needed; (ii) for some families of graphs, $\mu$ is a lower bound on the number of tests; and (iii) the lower bound cannot be described using only $\mu$ and $H(X)$. In particular, we provide examples where the tight lower bound is greater than the entropy but less than $\mu$. We further study the performance of the proposed algorithm for random uniform hypergraphs (where each hyperedge of size $d$ is present with probability $r$) and shed light on the optimality of the algorithm as a function of the hypergraph density $r$.
  • We introduce a semi-non-adaptive group testing model, a model in between adaptive and non-adaptive group testing, in which tests are performed sequentially, allowing the algorithm to halt when it finds the infected set. But, the model is semi-non-adaptive in that tests are designed independent of the results of previous tests. For this model, and assuming concentration of the infection size to be around its mean $\mu$, we provide an algorithm that requires ${O}(\mu H(X)+\mu\log n)$ tests. We also highlight challenges in designing non-adaptive algorithms. Specifically, we show that, unlike classic independent group testing, there can be a significant gap in the number of tests needed in adaptive vs non-adaptive testing.
  • We extend our methods and results to noisy group testing (where test results are corrupted by a symmetric noise) in both adaptive and semi-non-adaptive settings.
2 Problem Formulation
2.1 Generalized Group Testing
We model a generalized group testing problem via hypergraphs. Consider the hypergraph $G=(V,E)$ where $V$ is the set of nodes and $E$ the set of hyperedges. Each hyperedge $e\in E$ is a subset of the nodes, $e\subseteq V$. We assume $|V|=n$. In a population of size n, each individual is represented by a node in this hypergraph and the hyperedges capture possible dependencies between the nodes’ states. A distribution $\mathcal{D}$ is assumed over the edges in $E$, and one edge $e\sim\mathcal{D}$ is sampled from this distribution to be infected. We use the terminologies “sampled” edge, “infected” edge, and “target” edge interchangeably. The probability that hyperedge $e$ is infected is denoted by $p(e)$. When a hyperedge is infected, all nodes $v\in e$ are infected and all other nodes $u\notin e$ are not infected. The goal is to perform group tests, with the minimum number of tests, until the infected edge is identified accurately.
Figure 1 show an example of a hypergraph $G=(V,E)$ where $V=\{v_{1},v_{2},v_{3},v_{4},v_{5}\}$ is the set of nodes, $E=\{e_{1}=\{v_{1},v_{2},v_{3}\},e_{2}=\{v_{1},v_{5}\},e_{3}=\{v_{4},v_{5}\}\}$ is the set of edges, and $p_{e_{1}}=0.3$, $p_{e_{2}}=0.2$, $p_{e_{3}}=0.5$ is the distribution over the edges. Note that $v_{2}$ and $v_{3}$ have a complete correlation and are always in the same state because each edge contains either both of them or none of them. Also, $v_{1}$ and $v_{4}$ are also completely correlated as they are always in opposite states because each edge contains exactly one of them.
To formulate the problem more precisely, let $X=(X_{1},X_{2},\ldots,X_{n})$ be the vector of the nodes’ states where $X_{i}=1$ if the $i$’th node is infected and $X_{i}=0$ otherwise. We denote the probability that node $i$ is infected by $p_{i}$. When an edge $e$ is sampled, we have $\forall v\in e:X_{v}=1$ and $\forall v\notin e:X_{v}=0$. Within this model, each node $v$ may belong to multiple edges and therefore, we have
$p_{v}=\mathbb{E}[X_{v}]=\sum_{e\in E:\ e\ni v}p(e).$
The expected number of infections, $\mu$, is thus given by
$\mu=\mathbb{E}[X_{1}+X_{2}+\cdots+X_{n}]=\sum_{v\in V}p_{v}.$ (1)
It is worthwhile to mention that in prior work such as [LCH+14], it is often assumed that $\mu\ll n$, based on the observation that otherwise the number of tests needed is $\Omega(n)$. Interestingly, this is not true in general when we have correlation, as captured in our model, and an example is given in Section 6.2.
We next define testing and recovery before formalizing our objective.
Testing.
A group test $T$ is defined by the subset of nodes $T\subseteq V$ that take part in that test. We denote the $i$’th test by $T_{i}$, $T_{i}\subseteq V$ and its result by $r_{i}$. When $r_{i}=1$, we say that the test is positive, and when $r_{i}=0$, we say that the test is negative.
Adaptive/non-adaptive/semi-non-adaptive testing. Testing can be adaptive or non-adaptive: In adaptive testing, each test $T_{i}$, $i=1,2,\ldots,L$, is a function of the prior tests $T_{1},\ldots,T_{i-1}$ as well as their respective results $r_{1},\ldots,r_{i-1}$. Note that the number of tests, $L$, is a random variable that can depend on the randomness in the algorithm, the randomly chosen target edge, as well as the randomness in the test results (if testing is adaptive and noisy). In non-adaptive testing, all the tests $T_{1},T_{2},\ldots,T_{L}$ are designed a-priorily and can be done in parallel. In other words, the design of the $T_{i}$’s are independent of their results.
We further introduce a semi-non-adaptive group testing model (SNAGT) that serves as an intermediate approach between non-adaptive and adaptive group testing. This model is relevant because it combines the advantages of both strategies: the design of individual tests is non-adaptive, meaning each test is planned without relying on the outcomes of previous tests. However, it introduces adaptiveness in the decision to stop testing, allowing the total number of tests, L, to be determined based on prior results. This is particularly practical in group testing scenarios where delays make it challenging to wait for each test result before deciding the next step. The ability to stop testing when necessary provides a flexible and efficient solution without the need for constant adjustments during the testing process. In Section 3.4, we present an algorithm for the adaptive testing. In Section 7, we examine non-adaptive and semi-non-adaptive algorithms, highlighting the challenges associated with each.
Noisy Testing. Test results can be noiseless or noisy. In the noiseless setting, we have $r_{i}=1$ (positive test result) iff there is at least one node in $T_{i}$ that is infected (i.e., belongs to the sampled edge) and we have $r_{i}=0$ otherwise. In the noisy setting, results are flipped with probability $\delta$, i.e. when $T_{i}$ contains a node from the target edge, $r_{i}=1$ with probability $1-\delta$ and $r_{i}=0$ with probability $\delta$. Similarly, when $T_{i}$ does not contain any node from the target edge, $r_{i}=0$ with probability $1-\delta$ and $r_{i}=1$ with probability $\delta$. In other words, we assume that the error is symmetric. The symmetric noisy model is studied in the literature for independent group testing in works such as [Sca18, SC18, CCJS11, MS98, CJBJ13]. In Section 8, we study this noisy testing for group testing with correlation.
Recovery.
Given the test sequence $T_{1},\ldots,T_{L}$ and the respective results $r_{1},\ldots,r_{L}$, the nodes’ states are estimated as $\hat{X}=(\hat{X}_{1},\hat{X}_{2},\ldots,\hat{X}_{n})$ where $\hat{X}_{i}$ is the estimated state of node $i$, $i=1,\ldots,n$. The probability of error is then defined as $P_{e}=P(X\neq\hat{X})$, where the randomness is over $\mathcal{D}$ (probability distribution over the edges) and possible randomness of the testing design.
Objective.
In this work, we aim to devise algorithms that, in expectation, use the least number of tests (minimize $\mathbb{E}[L]$) and recover the sampled edge with probability of error $P_{e}$. The error probability $P_{e}$ can be zero, a value that goes to zero as $n$ goes to infinity, or a fixed goal error.
Note that the above model is equivalent to arbitrary correlation among nodes. Consider an arbitrary distribution $\mathcal{P}(S)$, $S\subset V=\{1,2,\ldots,n\}$, defined over the power set of a population size $n$, encoding the probability that exactly subset $S$ of the population is infected. Then, there is a corresponding hypergraph $G$ where an edge $e=S\subseteq V$ being the target edge is equivalent to set $S$ being infected, and $\mathcal{D}(e)=\mathcal{P}(S)$. It is important to note that the set of hyperedges can also be a proper subset of the power set, depending on the distribution $\mathcal{D}$, so we have $|\text{support}(\mathcal{D})|=|E|$.
2.2 Prior Models as Special Cases
We next discuss how different statistical models for infection of nodes in group testing can be considered as a special case of the above model. In particular, we find the probability mass function $\mathcal{D}(S)$ for all possible hyperedges $\mathcal{S}\subseteq V$.
First consider independent group testing [LCH+14]. In this model, we have $\mathcal{D}(X)=\Pi_{i}[X_{i}p_{v_{i}}+(1-X_{i})(1-p_{v_{i}})]$. Going beyond independent group testing, we consider models where correlation is modeled by edge faulty graphs [AU23, NKC+22]. [NKC+22] considers a simple graph $H=(V_{H},E_{H})$ where each (simple) edge is sampled in $H$ randomly with the same probability and the nodes that end up in the same component are assumed to have the same state, independently of the nodes in other components. Translating to the hypergraph model, for a binary vector $X=(X_{1},\ldots,X_{n})$, let $S=\{i\mid X_{i}=1\}$. Then $\mathcal{D}(S)$ is the probability that the nodes in $S$ (and only those) are infected in $H$. One can compute $\mathcal{D}(S)$ by taking a sum over the probability of graphs in which $S$ is disconnected from the rest of the graph, all the components formed in $S$ are infected and all the components formed in $V\setminus S$ are not infected. This gives us the equivalent hypergraph model for [NKC+22]. For more details, please refer to Appendix A. In a related work in [AU23], a similar model is considered, but with the restriction that only one component can be infected, and the infection probabilities across edges may vary. Again, we can build the equivalent hypergraph model by computing the probability mass function $\mathcal{D}(S)$ for every hyperedge $S$. This is done by taking the sum over probabilities of graphs where $S$ is a connected component after dropping the edges and is the only infected component. For more details, please refer to Appendix A.
In [ACÖ23], the authors consider $m$ communities each of size $k$. Initially, each node is a “seed” with probability $q$. Then the seeds infect nodes in the same community with probability $q_{1}$ and outside of the community with probability $q_{2}<q_{1}$. The probability that exactly one set $S$ is infected is the probability that no one is seed in $V\setminus S$ (where $V$ is the union of all communities) and set $S$ becomes infected with $i,1\leq i\leq|S|$ seeds. For more details, please refer to Appendix A.
In [NSG+21a], the authors consider $F$ families where each family is infected with probability $q$ independently. Then each node in an infected family $F_{j}$ is infected with probability $p_{j}$ independent of the rest, and the non-infected families have no infected nodes. The probability that a subset $S$ of the individuals are infected is that first, their family is infected (with probability $q$) and second, if $i\in S$ is in $j$’th group, $i$ is infected (with probability $p_{j}$). Now if $S=\{v_{1},\ldots,v_{k}\}$, and $I_{j}$ is the number of nodes of $F_{j}$ in set $S$, then the probability that only set $S$ becomes infected is
$\mathcal{D}(S)=\prod_{\begin{subarray}{c}j:I_{j}>0\end{subarray}}q(1-p_{j})^{|F_{j}|-I_{j}}(p_{j})^{I_{j}}\cdot\prod_{\begin{subarray}{c}j:I_{j}=0\end{subarray}}(1-q+q(1-p_{j})^{F_{j}}).$
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$e_{1}\text{, $p_{e_{1}}=0.3$}$$e_{2}\text{, $p_{e_{2}}=0.2$}$$e_{3}\text{, $p_{e_{3}}=0.5$}$
Figure 1: A hypergraph with $V=\{v_{1},v_{2},v_{3},v_{4},v_{5}\}$ and $E=\{\{1,2,3\},\{1,5\},\{4,5\}\}$ where $p_{\{1,2,3\}}=0.3$, $p_{\{1,5\}}=0.2$, and $p_{\{4,5\}}=0.5$.
2.3 Preliminaries
In this section, we provide the tools and basic ideas that we will use to describe and analyze our proposed group testing algorithm. We first establish a lower bound on the number of tests needed. Our result is inspired by [LCH+14] where almost matching lower and upper bounds are proved for adaptive and non-adaptive independent group testing. Following a similar argument, we prove a lower bound for the hypergraph problem. Specifically, in [LCH+14] they proved the following result:
Theorem 2.1.
[LCH+14] For the case where nodes are independent, i.e. $\mathcal{D}(X)=\Pi_{i}[X_{i}p_{v_{i}}+(1-X_{i})(1-p_{v_{i}})]$, for any algorithm that recovers the infection set and performs $L$ tests with probability $1-\epsilon$ we have
$L\geq(1-\epsilon)H(X).$
On the other hand, there is an adaptive algorithm that finds the infected set with a probability that goes to 1 as $n\rightarrow\infty$ using $L\leq O(\mu+H(X))$ tests where $H(X)=\sum_{v\in V}-p_{v}\log p_{v}$ and $\mu$ is given by (1).
The proof provided for the $(1-\epsilon)H(X)$ lower bound is quite general and can be extended to our work as following:
Theorem 2.2.
For any algorithm that recovers the sampled edge $e^{*}$ with probability at least $1-\epsilon$ using $T$ tests, we have
$L\geq(1-\epsilon)H(X)$ (2)
where $H(X)=\sum_{e\in E}-p(e)\log p(e)$.
Remark 2.3.
In contrast to the result of Theorem 2.1 where $H(X)$ characterizes the minimum number of tests (up to an additive factor), $H(X)$ is not necessarily tight in our setting. We discuss this more later in Example 6.1 where $H(X)=O(\log n)$ but $\Omega(n)$ tests are needed.
3 Adaptive Testing: An Algorithmic Approach
3.1 Challenges
Most classic adaptive algorithms for independent group testing are, in essence, built on generalized binary-splitting which greedily chooses the test that most evenly splits the candidate nodes or equivalently chooses the test with the maximal information gain. In probabilistic group testing (with independent nodes) [LCH+14], this involves choosing subsets of nodes to test such that the probability of a subset containing an infected node is close to $1/2$ (meaning that the test provides close to one bit of information). A negative test result allows a subset to be ruled out. Candidate subsets that test positive are, however, partitioned into two subsets (with roughly equal probability mass) for further testing. With such a design, the posterior probability of the sets testing positive, given the sequence of previous test results, is close to $1/2$. To follow this paradigm, we face three challenges.
First and foremost, it is important to note that treating each node individually and disregarding the underlying correlation is not efficient in terms of the number of tests needed and we aim to propose group testing strategies that exploit the correlation.
Under this model, even verifying whether a specific edge $e$ is the target edge or not requires many tests. As a matter of fact, it may need a number of tests that are of the order of its size. For example, consider the hypergraph in Figure 2 with $k=4$ nodes and consider $e_{1}$ as the suspected edge set with size $k-1$. To verify whether $e_{1}$ is the target edge, we cannot treat it as a whole and conclusively determine if $e_{1}$ is the actual edge or not based on a single test: If the test is positive, it can be $e_{1}$ that is realized or any other edge. In fact, any test of size greater than one is positive. Indeed, it takes $\Omega(k)$ tests to determine if $e_{1}$ is the target edge. This is because if $e_{1}$ is not the target edge, only a single node is not infected, and detecting the negative node will take $\Omega(k)$ individual tests.
$v_{1}$$v_{2}$$v_{3}$$v_{4}$
Figure 2: A graph with 4 nodes and 4 edges. Each edge contains 3 nodes, for example, $e_{2}=\{v_{1},v_{3},v_{4}\}$
Now consider the classical idea of greedily testing subsets and ruling out those subsets that test negative. If one wishes to utilize the correlation that is inherent in our model, it is not straightforward how the group tests should be designed: (i) Suppose that you have chosen a subset $S$ of nodes to test. If this test is negative, not only all nodes in $S$ are negative but also all those edges that contain a node in $S$ are not the target infected edge and could be ruled out. (ii) If the test is positive, we cannot conclude that the nodes that belong to $V\backslash S$ are negative. As a matter of fact, all can still remain valid candidates if they share an edge with an element of $S$. In other words, we can not conclude that the target infected edge is a subset of the nodes of set $S$. The design of the tests is thus nontrivial.
In the following, we demonstrate in Lemma 4.2 how a test can rule out a set of edges. Using this lemma, and updating the posterior probabilities of infections after each test, we design a greedy algorithm to sequentially rule out negative edge sets.
3.2 Some Useful Definitions
In order to design the tests sequentially, we utilize the posterior probability of nodes and edges of the hypergraph being infected given the previous tests and their results. In particular, suppose tests $T_{1},T_{2},\ldots,T_{k}$ are previously performed and the results are $r_{1},r_{2},\ldots,r_{k}$. We show the posterior probability of edge $e$ with
$q_{e\mid\{(T_{1},r_{1}),(T_{2},r_{2}),\ldots,(T_{k},r_{k})\}}.$
Note that $q_{e|\{\}}=p(e)$. Posterior probability of $v\in V$ being infected is similarly defined by
$q_{v\mid\{(T_{1},r_{1}),(T_{2},r_{2}),\ldots,(T_{k},r_{k})\}}=\sum_{e\in E:\ e\ni v}q_{e\mid\{(T_{1},r_{1}),(T_{2},r_{2}),\ldots,(T_{k},r_{k})\}}.$ (3)
When it is clear from the context, we drop $\{(T_{1},r_{1}),(T_{2},r_{2}),\ldots,(T_{k},r_{k})\}$ and use $q_{e}$ and $q_{v}$.
To connect edge probabilities to the probability that a test becomes positive, we need the following definition.
Definition 3.1.
For a subset $S\subseteq V$, the edge set of $S$ is defined as $E(S)=\{e\mid e\in E\wedge\forall v\in e:v\in S\}$. The weight of the set $S$ after tests $\{(T_{1},r_{1}),(T_{2},r_{2}),\ldots,(T_{k},r_{k})\}$ is defined as
$w(S|\{(T_{1},r_{1}),(T_{2},r_{2}),\ldots,(T_{k},r_{k})\})=\sum_{e\in E(S)}q_{e|\{(T_{1},r_{1}),(T_{2},r_{2}),\ldots,(T_{k},r_{k})\}}.$ (4)
When clear from the context, we drop $\{(T_{1},r_{1}),(T_{2},r_{2}),\ldots,(T_{k},r_{k})\}$ and simply write $w(S)$.
After performing $k$ tests and seeing the results, the expected number of infections would be updated according to the posterior probabilities discussed above. We denote the expected number of infected nodes in the posterior probability space by:
$\mu_{\{(T_{1},r_{1}),(T_{2},r_{2}),\ldots,(T_{k},r_{K})\}}=\sum_{v}q_{v\mid{\{(T_{1},r_{1}),(T_{2},r_{2}),\ldots,(T_{k},r_{k})\}}}$ (5)
We drop ${\{(T_{1},r_{1}),(T_{2},r_{2}),\ldots,(T_{k},r_{k})\}}$ when it is clear from the context, and show it with $\tilde{\mu}$ to avoid confusion with the original $\mu$ before performing any test.
Throughout this work, when a new test $(T_{i},r_{i})$ is done, we compute $q_{e|(T_{1},r_{1}),\ldots,(T_{i},r_{i})}$ from the last posterior $q_{e|(T_{1},r_{1}),\ldots,(T_{i-1},r_{i-1})}$. In other words, we initially start with $\mathcal{D}_{0}=\mathcal{D}$ and after observing the results of the $i$’th test, compute $\mathcal{D}_{i}$ from $\mathcal{D}_{i-1}$ based on $(T_{i},r_{i})$ where $\mathcal{D}_{i}=\{q_{e|(T_{1},r_{1}),\ldots,(T_{i},r_{i})}\}_{e\in E}$. In Lemma 4.3, we prove that computing the posterior probability entails “removing some edges” from the hypergraph and “scaling the distribution” according to the following definition:
Definition 3.2.
Let $\mathcal{D}$ be a distribution over a set $E$. Removing $e\in E$ from $\mathcal{D}$ entails setting $\mathcal{D}(e)=0$ and re-normalizing the distribution by a factor $c>1$, $\mathcal{D}(e)\leftarrow c\mathcal{D}(e)$, such that $\sum_{e\in E}\mathcal{D}(e)=1$.
[…]
3.3 Overview and Ideas
Figure 3: An illustration of $E(S)$. Here, $E(S)=\{e_{1},e_{2}\}$ but $e_{3}\notin E(S)$ as one of its endpoints is outside of $S$. Hence, $w(S)=p_{e_{1}}+p_{e_{2}}.$ Test $V\setminus S$ is positive iff $e^{*}=e_{3}$ or $e^{*}=e_{4}$.
The ideas behind our algorithm can be summarized in three points:
  • Devise a testing mechanism in which by testing a subset of the nodes, we can rule out edge sets that are not compatible with the result.
  • Design the test so that, independent of the test result, a constant fraction of the mass is ruled out. Here, by a constant, we mean a mass that is independent of all the problem’s parameters such $G$ and $\mathcal{D}$. We argue that a most informative test can provide at most one bit of information by ruling out edge sets that constitute half of the probability mass. By removing a constant fraction of the mass, we make sure that our tests are near-optimal.
  • Updating posterior probabilities of infections for the nodes and edges based on the prior test results. It turns out that this is equivalent to removing edges that are ruled out from the hypergraph, and scaling the remaining edge probabilities uniformly up so they sum up to 1. We perform such posterior updates sequentially after each group test. If at every step we can do this until we reach a single edge (with probability 1), then we would have done $O(H(X))$ tests in expectation (Theorem 4.8) which is optimal by the matching lower bound in Theorem 2.2. These types of tests are “informative”, as they shrink the search space significantly.
If we can not find such a test, we observe that some nodes have a high probability of infection and other nodes are likely to be negative. (Proposition 4.4). By testing the “unlikely” nodes, we either rule them all out if the test is negative and are left with nodes that have a high probability of infection, or will have an informative test if the test is positive as we can rule out a high mass of the edges.
Suppose a set $S$ with $w(S)=1/2$ is given and we test $V\setminus S$. If the test is positive, it means that the sampled edge $e^{*}$ is not in $E(S)$ which contains half the probability mass. If the test is negative, then $e^{*}$ is in $E(S)$ which rules out $E\setminus E(S)$. So in either case, we gain one bit of information. Figure 3 illustrates $S$, $E(S)$ and $V\backslash S$. Since finding a set that has $w(S)=1/2$ might not be possible, we relax this condition. We aim to find a set $S$ so that $w(S)$ is between two constants $c=1/2-\delta$ and $1-c=1/2+\delta$, let’s say between $0.05$ and $0.95$. Then by testing $V\setminus S$, we can understand if the sampled edge is in $E(S)$ or not. Similar to the case of $c=1/2$, either $E(S)$ or $W\setminus E(S)$ will be ruled out, which has a mass of at least $0.05$. There is a nuance in the definition of $E(S)$, which contains all the edges that are completely inside of $S$. If we had included the edges that have at least one node in $S$ as $E(S)$, then we couldn’t have ruled out the edges as we did.
We next describe how to find $S\in V$ with .$95.05\leq w(S)\leq.$95. We design a greedy-like algorithm that finds $S$ iteratively by removing the nodes from the working set. The algorithm starts with $S=V$. Initially, $w(S)=1$. Then the algorithm drops the nodes in $S$ one by one until it finds $0.05\leq w(S)\leq 0.95$ (at which point the algorithm performs a test on $V\setminus S$ and based on the result updates the posterior distribution).
What happens if such an $S$ does not exist? In this case weight of every subset of nodes is either more than $0.95$ or less than $0.05$, and removing any node from any subset whose weight exceeds $0.95$ reduces the weight of the subset to below $0.05$. This can for example happen when each node whose status is yet to be known definitively either is infected with a very high probability (exceeding $0.9$) or with a very low probability (a set of all such nodes will test positive with probability at most $0.05$). What we propose in this stage is the following: in a single test, the algorithm can test all the low-probability nodes, and if it is positive, we rule out $0.95$ of the mass, which is an unexpected gain, and continue with the algorithm. If it is negative, we end up only with nodes that all have high probabilities. Then the algorithm tests nodes individually.
We refer to the first stage of the algorithm when an informative test can be found based on $S$, $c<w(S)<1-c$, as Stage 1 of the algorithm and the second stage of the algorithm for which no $S$, $c<w(S)<1-c$, can be found as Stage 2 of the algorithm.
In Corollary 4.6, we show that in Stage 1, where each test removes at least $c$ fraction of the mass, the algorithm performs at most $O(H(X))$ tests in expectation which is near-optimal. In Stage 2, where high probability nodes are tested individually, we show in Observation 4.7 that the number of tests depends on the expected number of infections at that point, $\tilde{\mu}$. This expectation is obtained by the posterior probability after performing all the previous tests. We prove bounds on $\tilde{\mu}$ under assumptions on the size of the edges. Specifically, when the size of the edges is concentrated, we get a bound in the form of $O(\mu+H(X))$, which is stated in Theorem 4.8. Another way to bound the number of tests in Stage 2 is to first drop the edges that are much larger than $\mu$, which have a low mass, and then run the algorithm. In this case, we can guarantee that $\tilde{\mu}$ is not much larger than $\mu$, and the guarantee for this approach is shown in Theorem 4.10.
3.4 The Proposed Algorithm
1 Input: Graph $G=(V,E)$, distribution $\mathcal{D}=\mathcal{D}$ over $E$, $c<1/2$, and the set $W$ containing negative nodes.
2 Output: Target edge $e^{*}$.
3 Initialize set $S=V\setminus W$.
4 Compute $w(S)$ by (4) using the (updated) posterior distribution $\mathcal{D}=\{q_{e}\}_{e\in E}$.
5 If $c\leq w(S)\leq 1-c$, test $V\setminus S$:
6     If the test is positive, remove $E(S)$ from $\mathcal{D}$ by def 3.2 to update the posterior distribution.
7     If the test is negative, remove $E\setminus E(S)$ from $\mathcal{D}$ by def 3.2 to update the posterior distribution. Set $W=W\cup(V\setminus S)$.
8     Return to line 1.
9 Else:
10     If there is an edge with $q_{e}=1$, return $e$.
11     If $\exists v:c\leq w(S\setminus v)\leq 1-c$, remove $v$ from $S$ and go to line 1.
12     If $\exists v:1-c<w(S\setminus v)$, remove $v$ from $S$ and go to line 1.
13     Otherwise ($w(S)>1-c$ and $\forall v:w(S\setminus v)<c$), test $V\setminus S$.
14        If the test is positive, remove $E(S)$ from $\mathcal{D}$ by def 3.2 to update the posterior distribution and go to line 1.
15        If the test is negative, test every node $v\in S$ individually with $10<q_{v}<$ 1, update the posterior after every test and return the nodes with $q_{v}=1$.
Algorithm 1 Adaptive Algorithm for General Group Testing
The proposed algorithm is presented in Algorithm  1. The algorithm undertakes two broad tasks repeatedly
  1. 1.
    Determine the set of nodes it tests, denoted as $V\setminus S$, (it identifies a set $S$ whose complement it tests);
  2. 2.
    If the set tests positive, only the edges which intersect with the tested set is retained, the rest are removed, if the set tests negative, all edges containing at least $1$ node in the set are removed, the tested set is removed from $V$ and added to set $W$, a set of nodes which are known to be negative; the posterior probability of each surviving edge is updated and $w(S)$ is calculated for each $S$ using (4) and the updated posteriors.
We now describe how these broad tasks are executed. In Stage 1 of the algorithm which corresponds to lines 1-1, the algorithm greedily finds a set $S$ so that $c\leq w(S)\leq 1-c$. This is captured in lines 1 and 1. When the set $S$ is found, $V\setminus S$ is tested in line 1 and the posterior is updated accordingly. Such a test helps rule out edges with total weight of at least $c$ are ruled out of contention for being target edge regardless of the outcome of the test. If the algorithm can not find a set $c\leq w(S)\leq 1-c$ at this stage, it tests $V\setminus S$ in line 1 which has a high probability of being negative. If at Line 1, the test result is positive (unexpectedly), the algorithm rules out $1-c$ fraction of the mass and gets back to line 1. Stage 2 of the algorithm is outlined in line 1, where it tests uncertain high probability nodes individually. Note that at this point, all the nodes outside of $S$ are negative.
We now provide details of the sequence followed in the algorithm. At the start of the algorithm, typically the set of nodes that are known to be negative, $W$, is empty. Thus in line 1, the set $S$ is initialized as $V\setminus W=V.$ Clearly in line 1 $w(S)=1$, and $S$ does not satisfy the condition in line 1. The algorithm would now move to line line 1 (typically the condition in line 1 would also fail as the target edge will not be clear at this point). In line 1, the algorithm identifies individual nodes which are tested in the loop starting from line 1, and testing these individual nodes rules out edges with total weight at least $c$ regardless of the outcome of the test. After each test, posteriori probabilities are modified and the graph is modified as in (2) (this modification happens in lines 1,1 and in line 1). The above sequence is repeated until there does not exist individual nodes such that by removing them $S$ satisfies the condition in line 1. Note that $w(S)=1$ each time the algorithm reaches lines 1. The tests are all individual while this duration lasts.
Once the duration described in the previous paragraph ends the algorithm moves to line 1. Note that $w(S)=1$ the first time that the algorithm reaches line 1 under the then posterior probabilities. When the algorithm reaches line line 1, $S$ consists only of nodes that are positive either with very high probability (such that if any one of those is removed, $w(S)$ falls below $c$ from $1$), or with very low probability (such that $w(S)$ exceeds $1-c$ even when any of those is removed from $S$). In line 1, the low probability nodes are identified one at a time, and once such a low probability node is identified, line 1 determines if now there exists a node whose removal ensures that $w(S)$ satisfies the condition in line 1. If there exists such a node, in line 1 it is tested together with the low probability node identified in line line 1. If there does not exist such a node, then the algorithm gets back to line 1 again and removes another low probability node and the same process is repeated. If no node is found in line 1, then all low probability nodes are tested together in line 1. Thus, tests are no longer individual once the algorithm reaches line 1. When a group of nodes is tested, either in line 1 or line 1, the graph is modified as in (2). If the test in line line 1 is negative, then all tested nodes are removed from $V$, and the remaining nodes are all positive with high probability and are tested individually in line line 1. If the test is positive, then only edges consisting of at least $1$ node in the set are retained, and the algorithm returns to line 1 and the sequence is repeated.
Before we start analyzing, we provide some details to get a better intuition of the algorithm and present two example. If an individual node (set) is tested negative, since it is removed from $V$, it is never tested again. If an individual node (set) is tested positive, only edges intersecting that node (set) remain in the graph, thus the weight of the complement of that node (set) becomes 00. Thus, the complement set will never have weight in the interval $[c,1-c]$ and is therefore not tested again.
Example 3.3.
Let’s run Algorithm 1 on the graph in Figure 1 with $c=0.1$. Initially, $S=\{v_{1},v_{2},v_{3},v_{4},v_{5}\}$ and $w(S)=1$. By removing $v_{1}$ from $S$, $E(S\setminus v_{1})$ contains only one edge $e_{3}=\{4,5\}$, hence $c<w(S\setminus v_{1})=0.5<1-c$ and the algorithm tests $v_{1}$. If it is negative, then it rules out $e_{1}$ and $e_{2}$ and after scaling the edges, $q_{e_{3}}=1$ and the algorithm returns it. If it is positive, it removes $e_{3}$ and scales the other edges by 2 so $q_{e_{1}}=0.6$ and $q_{e_{2}}=0.4$. Again $S$ would be initialized to contain all the nodes, and after removing $v_{2}$, $w(S\setminus v_{2})=0.4$ so the algorithm test $v_{2}$ to eliminate one of $e_{1}$ or $e_{2}$. The expected number of tests here is 1.5.
In the following, we give a more abstract and larger example.
Example 3.4.
To give another illustrating example, suppose there are $k$ islands where each island has $m$ nodes, so $n=mk$. Nodes on the same island are perfectly correlated, ie, they always have the same state, as such every hyper edge always contains all or none of any given island (Figure 4, $k=4,m=3$). Thus, there are $2^{k}$ hyperedges at most; note that in this case the number of hyper edges can be significantly smaller than than the size of the power set of nodes, $2^{mk}$. Consider a setting with three types of islands: low, moderate and high, characterized by their probability of infection. For all nodes in high islands, $p_{i}\approx 1$ (e.g. island 4 in Figure 4). In moderate islands (e.g. islands 2,3), $p_{i}\approx\frac{1}{2}$ and in low islands, e.g. island 1, $p_{i}\approx\epsilon$, where $\epsilon$ is a small, positive number. One can verify that Algorithm 1 first picks a candidate from one of the moderate islands in line 1 and tests it individually. If it is positive, posterior probabilities of all nodes in its island is updated to $1$, otherwise to 00, through an update of posterior probabilities of all the edges (line 1 or 1). Accordingly, no other node from this island is identified for testing in line 1. Subsequently, similarly one node from each moderate island is identified for testing in line 1, and posterior probabilities are updated which eliminates the need to test other nodes from the same island in line 1. At the end of this process, all islands are high or low. Then, the algorithm gathers all the nodes from all low islands, in line 1 and tests them at once in line 1. With high probability, the test is negative, and finally, the algorithm tests nodes in high islands individually, except those whose posterior probability is already updated to $1$, updates the posterior probabilities of other nodes after each test, and returns all the nodes with $q_{v}=1$. Note that algorithm operates such that only one node is tested from each island.
Figure 4: An example with $k=4$ islands. Each edge contains either all nodes in an island or none of it. Three edges are given in red. Nodes in island 1 are infected with $p_{1}\approx 0$, nodes in islands 2 and 3 are infected with probability $p_{2}=p_{3}\approx 1/2$, and nodes in island 4 are infected with high probability $p_{4}\approx 1$.
4 Analysis
In this section, we bound the expected number of tests performed by Algorithm 1. The main result is built on the following key lemma. Here, $H(X)$ is the entropy of the nodes and $\tilde{\mu}$ is the expected number of infections in line 1 (computed with the posterior probabilities).
Lemma 4.1.
Algorithm 1 finds $e^{*}$ with $\mathbb{E}[L_{1}+L_{2}]\leq\frac{1}{\log\frac{1}{(1-c)}}H(X)+\frac{\mathbb{E}[{\tilde{\mu}}]}{1-2c}$ tests in expectation.
The roadmap of our proof for Lemma 4.1 is as follows. We first prove that removing $E(S)$ (resp. $E\backslash E(S)$) from $\mathcal{D}$ (see Definition 3.2) upon obtaining a positive (resp. negative) test result for $V\backslash S$ provides the true posterior distribution. This is established in Lemma 4.3. We next bound the number of tests $L_{1}$ done in Stage 1 of the algorithm by $\frac{1}{\log\frac{1}{(1-c)}}H(X)$, in expectation, and then show that the number of individual tests $L_{2}$ done in Stage 2 is bounded by $\frac{{\tilde{\mu}}}{1-2c}$.
Posterior Update.
We first need the following lemma to justify the correct operation of Algorithm 1 in removing edges after each test. In particular, the target edge $e^{*}$ is never removed and each test indicates whether the target edge is in $E(S)$ or $E\setminus E(S)$.
Lemma 4.2.
The result of a test $V\setminus S$ is positive iff $e^{*}\in{E\setminus E(S)}$.
Proof.
A test $V\setminus S$ is positive if at least one of the nodes in $V\setminus S$ is positive, which means $e^{*}$ contains a node in $V\setminus S$. But by definition, $E(S)$ is the set of edges with all nodes in $S$, hence $e^{*}\notin E(S)$ and $e^{*}\in E\setminus E(S)$. Similarly, if $e^{*}\in E\setminus E(S)$, it contains at least one node in $V\setminus S$ and hence the result is positive. ∎
Building on Lemma 4.2, we further prove that the update rule utilized in Algorithm 1 computes the posterior probabilities given all the previous tests.
Lemma 4.3.
Consider a distribution $\mathcal{D}$ over the edge set $E$ and a test $T=V\setminus S$ with result $r$. If $r=1$, the posterior probability is obtained by removing $E(S)$ according to Definition 3.2. If $r=0$, the posterior probability is obtained by removing $E\setminus E(S)$ according to Definition 3.2.
Proof.
For simplicity, consider the update after the first test and $r=1$. Then the posterior is
$q_{e_{1}\mid(T_{1}=S\setminus V,r_{1}=1)}$ $=Pr(e_{1}=e^{*}|T_{1}=S\setminus V,r_{1}=1)$ (6)
[…]
$=Pr(e_{1}=e^{*},e^{*}\in E\setminus E(S))/P(e^{*}\in E\setminus E(S))$ (8)
=$=\begin{cases}0,&\text{ if }e_{1}\notin E\setminus E(S)\\ Pr(e_{1}=e^{*})/(\sum_{e\in E\setminus E(S)}p(e))&\text{otherwise}\end{cases}$ (9)
The case for $r_{1}=0$, a negative test, can be argued similarly. The main difference in this case is that we will prove the reverse of lemma 4.2, ie, $r_{1}=0$ iff $e^{*}\in E(S)$. Now, $r_{1}=0$ needs to be substituted with $r_{1}=1$, and $E(S)$ with $E\setminus E(S)$ in the above equations, and otherwise the same result can be obtained.
For tests $i,i>1$, Lemma 4.2 holds for $r_{i}=1$, and considering $E$ updated after $i-1$ tests. Equations (6) to (9) follow with the conditioning updated to include results of all tests and using posterior probabilities $q_{e}$ instead of $p(e)$. The same equations can be derived for $r_{i}=0$ as in the previous paragraph. ∎
Informative Tests in Stage 1.
The above lemma implies that if both $E(S)$ and $E\setminus E(S)$ have a moderate mass (i.e., $w(S)\geq c$ and $1-w(S)\geq c$), regardless of the result, the test is informative. This is quantified next for Stage 1 of the algorithm. We first give a structure on probabilities of nodes in line 1, where the algorithm fails to find a set $S$ so that $c\leq w(S)\leq 1-c$.
Proposition 4.4.
When Algorithm 1 reaches  1, we have (i) $w(S)>1-c$, (ii) $Pr(\exists v\in V\setminus S:v\in e^{*})<c$ where $e^{*}$ is the sampled edge, and (iii) $\forall v\in S:q_{v}>1-2c$.
Proof.
To see (i) note that at line 1 $w(S)=1>1-c$ and, as we proceed by removing nodes from $S$ to meet the criteria of Line 1, it never gets met and therefore when the algorithm reaches Line 1, we have $w(S)>1-c$. To prove (ii), we compute the probability that at least one node $v\in V\setminus S$ is positive. By Lemma 4.2, this is equal to the probability that $e^{*}\in E\setminus E(S)$, i.e., $\sum_{e\in E\setminus E(S)}q_{e}=1-w(S)$ which is at most $c$ by (i). Finally, to prove (iii), we note that $v\in S$, $q_{v}\geq w(S)-w(S\setminus v)$, because some of edges that contain $v$ have nodes outside of $S$. But the algorithm has passed line 1, so $w(S)>1-c$ and $w(S\setminus v)<c$ for any node $v$. Therefore, $q_{v}\geq w(S)-w(S\setminus v)>1-2c$.
The following observation shows that the tests in the first stage are “informative”.
Observation 4.5.
For Algorithm 1, in Stage 1, for every update of $\mathcal{D}$ (lines 1, 1,1), at least $c$ fraction of the mass is removed.
Proof.
If the algorithm scale $\mathcal{D}$ at lines 1 or 1, by lemma 4.2, we either remove $E(S)$ or $E\setminus E(S)$. But $c\leq w(S)\leq 1-c$, so in either case we remove $c$ fraction of the mass.
If the algorithm scales $\mathcal{D}$ at line 1, it means the test had been positive. By Proposition 4.4, $w(S)=1-Pr(v\in V\setminus S\cap v\in e^{*})>1-c>c$ because $1-c>1/2>c$, meaning at least $c$ mass have been removed. So for any update, at least $c$ fraction is removed. ∎
Now we can bind the number of tests $L_{1}$ performed in the first stage of the algorithm using the fact that each test before line 1 is informative.
Corollary 4.6.
Before line 1, the number of tests $L_{1}$ that Algorithm 1 performs is bounded in expectation as follows:
$\mathbb{E}[L_{1}]\leq\frac{1}{\log\ \frac{1}{(1-c)}}H(X).$
Proof.
By Observation 4.5, we know that every time the posterior $\mathcal{D}$ is updated in Stage 1, at least a mass $c$ is removed and the remaining edge probabilities are scaled by at least a factor of $\frac{1}{1-c}$. Suppose that the target edge $e^{*}$ is selected with probability $p_{e^{*}}$. It takes at most $\log_{\frac{1}{1-c}}(1/p_{e^{*}})$ tests for the posterior on $e^{*}$ to become 1. So the number of tests before line 1 is at most $\log_{\frac{1}{1-c}}(1/p_{e^{*}})=\frac{\log 1/p_{e^{*}}}{\log(\frac{1}{1-c})}$, leading to the following bound on $\mathbb{E}[L_{1}]$:
$\mathbb{E}[L_{1}]<\sum_{e\in E}p(e)\log_{\frac{1}{1-c}}(p(e))\\ =\sum_{e\in E}p(e)\frac{\log 1/p(e)}{\log\frac{1}{1-c}}\\ =\frac{1}{\log\frac{1}{1-c}}\sum_{e\in E}p(e)\log 1/p(e)\\ =\frac{1}{\log\ \frac{1}{(1-c)}}H(X).$
Individual Testing in Stage 2.
Algorithm 1 stops performing in Stage 1 when the greedy process of finding $S$, $c\leq w(S)\leq 1-c$, fails in line 1, and the nodes out of $S$ tests negative in line 1. At this point, in Stage 2, we resort to individual testing in Line 1.
Using Proposition 4.4, we bind the expectation of $L_{2}$ based on the expectation at the “stopping time” $\tilde{\mu}$ in the following.
Lemma 4.7.
Let $\tilde{\mu}=\mu_{(\mathbf{T_{1},r_{1}),\ldots,(T_{L_{1}},r_{L_{1}})}}$ be the expected number of infected nodes at the end of line 1 where $\mathbf{(T_{i},r_{i})},1\leq i\leq L_{1}$ are random variables which indicate the tests and results performed. Then the algorithm performs at most $L_{2}\leq\frac{\tilde{\mu}}{1-2c}$ tests in line 1.
Proof.
Suppose test $\mathbf{(T_{i},r_{I})}=(T_{i},r_{i})$ is done. In other words, we condition $\tilde{\mu}$ to the case where tests are $(T_{i},r_{i})$’s. First recall that at any step of the Algorithm, $q_{v}$ stands for $q_{v\mid\{\{(T_{1},r_{1}),\ldots,(T_{L_{1}},r_{L_{1}})\}}$. Let $X_{v}$ be the indicator variable representing if node $v$ is infected or otherwise. Thus, $q_{v}=E[X_{v}\mid\{(T_{1},r_{1}),\ldots,(T_{L_{1}},r_{L_{1}})\}].$ and
$\tilde{\mu}\mid\{\mathbf{(T_{1},r_{1})}=(T_{1},r_{1}),\ldots,(\mathbf{T_{L_{1}},r_{L_{1}})}=(T_{L_{1}},r_{L_{1}})\}$
$\triangleq\mathbb{E}[\sum_{v}X_{v}\mid\{(T_{1},r_{1}),\ldots,(T_{L_{1}},r_{L_{1}})\}]$
$=\mathbb{E}[\sum_{v\in S}X_{v}\mid\{(T_{1},r_{1}),\ldots,(T_{L_{1}},r_{L_{1}})\}]$
$=\sum_{v\in S}\mathbb{E}[X_{v}\mid\{(T_{1},r_{1}),\ldots,(T_{L_{1}},r_{L_{1}})\}]$
$=\sum_{v\in S}q_{v}>|S|(1-2c)$
Recall that $S$ here is defined in Algorithm 1 when it reached Line 1. The first equality is true because in the previous line, $V\setminus S$ tested negative, so $X_{v}=0$ for $v\in V\setminus S.$ The second inequality is true because given the test results and which nodes have been tested thus far, the set $S$ is specified, thus the expectation can be moved inside the summation. The third equality follows from what $q_{v}$ stands for. The fourth inequality follows from Proposition 4.4, from which we know that for each $v\in S$, $q_{v}>1-2c$. But the algorithm performs $|S|$ tests at line 1, so $L_{2}=|S|$. Thus regardless of the test results, $L_{2}<\frac{\tilde{\mu}}{1-2c}.$
The above observation in conjunction with Corollary 4.6 proves Lemma 4.1.
Theorem 4.8 below is a direct implication of Lemma 4.2 when the size of the edges do not differ by much.
Theorem 4.8.
If $\forall e\in E:f_{1}(n)\leq|e|\leq f_{2}(n)$, then Algorithm 1 finds $e^{*}$ with $\frac{1}{\log\frac{1}{(1-c)}}H(X)+\frac{f_{2}(n)}{f_{1}(n)(1-2c)}\mu$ tests in expectation.
Proof.
Recall that $\mu$ is the average number of infections before performing any tests. As $f_{1}(n)\leq|e|$ for each edge $e$, we have $\mu\geq f_{1}(n)$. As $|e|\leq f_{2}(n)$, we have $\tilde{\mu}\leq f_{2}(n)$. Then the ratio $\frac{\tilde{\mu}}{\mu}\leq\frac{f_{2}(n)}{f_{1}(n)}$, hence $\mathbb{E}[\tilde{\mu}]\leq\frac{f_{2}(n)}{f_{1}(n)}\mu$. Replacing this in $\mathbb{E}[L_{1}+L_{2}]=\frac{1}{\log\frac{1}{(1-c)}}H(X)+\frac{\mathbb{E}[\tilde{\mu}]}{1-2c}$ from Lemma 4.1 completes the proof. ∎
Note that if most of the edges have size concentrated around $\mu$, we can remove very large or very small edges and incur a small probability of error, but get a better bound. This idea is summarized in the following corollary.
Corollary 4.9.
If there is an $E^{\prime}\subseteq E$ such that $\sum_{e\in E^{\prime}}p(e)=1-\epsilon$ and $\forall e\in E^{\prime}:f_{1}(n)\leq|e|\leq f_{2}(n)$, then Algorithm 1 finds $e^{*}$ with $\frac{1}{\log\frac{1}{(1-c)}}H(X)+\frac{f_{2}(n)}{f_{1}(n)(1-2c)}\mu$ tests in expectation with probability $1-\epsilon$.
While Theorem 4.8 imposes a constraint on the edge sizes and finds the infected set with probability 1, the constraint can be relaxed if a small probability of error ($\epsilon$, fixed or approaching zero) can be tolerated. We use Markov Lemma to bound the probability that $|e^{*}|$ exceeds $\mu/\epsilon$. By removing those edges for which $|e|>\mu/\epsilon$, we ensure $\tilde{\mu}<\mu/\epsilon$ and establish the following result.
Theorem 4.10.
There is an algorithm that performs $\mathbb{E}[L_{1}+L_{2}]\leq\frac{1}{\log\frac{1}{(1-c)}}H(X)+\frac{\mu}{\epsilon(1-2c)}$ tests in expectation and returns the correct edge $e^{*}$ with probability $1-\epsilon$.
Proof.
Let $z=|e^{*}|$. Then we know $\mathbb{E}[{z}]=\mu$ and by Markov inequality, $Pr[z>\mu/\epsilon]<\epsilon$, meaning that with probability $1-\epsilon$, the size of the infected set is at most $\mu/\epsilon$. Now we can modify Algorithm 1 to exploit this fact. The revised algorithm maintains the same execution as Algorithm 1, however, before initiation, it removes all the edges that have a size greater than $\mu/\epsilon$.
We know that with probability $1-\epsilon$, the realized edge is not removed and the largest edge has size at most $\mu/\epsilon$. Therefore, with probability $1-\epsilon$, the algorithm returns the correct $e^{*}$. The number of tests is argued as before, with the difference that the size of $\tilde{\mu}\leq\mu/\epsilon$ (because the largest edge has size at most $\mu/\epsilon$), so in the second round it performs at most $L_{2}\leq\frac{\mu}{\epsilon(1-2c)}$ tests and the proof is completed. ∎
In the above Theorem, there is a trade-off between the probability of error and the expected number of tests. With $\epsilon$ error, the second term of expectation increases with $1/\epsilon$. So if we set $\epsilon\rightarrow 0$ when $n\rightarrow\infty$, for example $\epsilon=1/\log n$, we incur a multiplicative factor of $\log n$ on the second term
Markov bound works for general distributions, but it might happen that the mass in $\mathcal{D}$ is concentrated around the mean. In this case, the dependency on $1/\epsilon$ might be slower. For example, if the concentration is exponential, i.e. $Pr(|e^{*}|>2\mu)<2^{-\Omega(n)}$, then with probability $1-2^{-\Omega(n)}$ the algorithm recovers $e^{*}$ and the expected number of tests is $\mathbb{E}[L_{1}+L_{2}]\leq\frac{1}{\log\frac{1}{(1-c)}}H(X)+\frac{2\mu}{(1-2c)}$. More precisely, we have the following theorem.
Corollary 4.11.
Suppose $e^{*}$ is sampled from the probability distribution $\mathcal{D}$ and we have $Pr(|e^{*}|>c\mu)=o(1)$, for a constant $c$. Then, with probability $1-o(1)$, Algorithm 1 returns the correct edge $e^{*}$ using $O(H(X)+\mu)$ tests with probability $1-o(1)$.
5 Recovering/Improving Prior Results
In this section, we consider the statistical models provided by prior related works and compare the results by the performance of Algorithm 1.
The work [LCH+14] considers independent group testing where node $i$ is infected with probability $p_{i}$ independent of the other nodes. The authors provide adaptive algorithms that use $O(\mu+H(X))$ tests on average and output the infected set with high probability. They also prove almost surely bounds on the number of tests. By applying Theorem 4.10 to this setting for a constant $c$, we obtain $O(H(X)+\frac{\mu}{\epsilon})$ as an upper bound on the expected number of tests where $\epsilon$ is the error probability. If $\mu=\sum_{i}p_{i}\gg 1$, which is a common assumption in the literature, then the number of infections is concentrated around $\mu$ and by Corollary 4.11, with probability $1-o(1)$, the algorithm in Theorem 4.10 recovers the infected set with $O(H(X)+\mu)$ tests, which recovers the adaptive upper bound result (Theorem 2 and Corollary 1) up to a constant provided in [LCH+14].
In [NSG+21a], as we discussed briefly in Section 2.2, the authors consider $F$ families where each family $j$ has size $M_{j}$, so $n=\sum_{j}M_{j}$. Each family $j$ is infected with probability $q$ and when a family is infected, all of its nodes are infected with probability $p_{j}$ independent of the rest. If a family is not infected, then all its nodes are not infected. They show that to have a constant probability of recovery, one needs $L\geq H(X)=Fh_{2}(q)+\sum_{j=1}^{F}qM_{j}h_{2}\left(p_{j}\right)-w_{j}h_{2}\left(\frac{1-q}{w_{j}}\right)$ where $w_{j}=1-q+q\left(1-p_{j}\right)^{M_{j}}$ and $h_{2}$ is the binary entropy. For the symmetric case where $\forall jp_{j}=p,M_{j}=M$ and hence $w_{j}=w$, the entropy simplifies to $H(X)=Fh_{2}(q)+qnh_{2}(p)-Fwh_{2}(\frac{1-q}{w})$. They provide an algorithm that in expectation performs at most $O(Fq(\log(F)+M)+nqp\cdot\log n)$.
Note that $\mu=nqp$, and by Theorem 4.10, with probability $1-\epsilon$ the target edge can be found using $O(Fh_{2}(q)+qnh_{2}(p)-Fwh_{2}(\frac{1-q}{w})+\frac{nqp}{\epsilon})$ tests for a constant $c$. If $F\gg 1$ and $M\gg 1$, then the expected number of infections is concentrated around its mean, and with probability $1-o(1)$ the algorithm succeeds using $O(Fh_{2}(q)+qnh_{2}(p)-Fwh_{2}(\frac{1-q}{w})+{nqp})$. This significantly improves the upper bound provided in [NSG+21a].
In [GLS22], the authors consider a combinatorial version of group testing where each edge has size $d$ and show that there is an algorithm that recovers $e^{*}$ with high probability using at most $O(\log|E|+d\log^{2}d)$ tests. If we put probability $1/|E|$ on each edge, then Theorem 4.8 tells us that we can recover the infected set with probability 1 using $O(H(X)+\mu)=O(\log|E|+d)$ tests in expectation, slightly improving the bound in [GLS22].
In [ACÖ23], the authors consider a stochastic block infection model (SBIM) where there are $m$ communities and each one has $k$ nodes, so $n=mk$. Each node is a seed with probability $p$ independently. Then every seed node infects nodes in the same community with probability $q_{1}$ and $q_{2}$ for the nodes outside of the community, $q_{1}>q_{2}$. The authors show lower and upper bounds on the number of tests. For the special case where $q_{2}=0$, they further show optimality of the proposed algorithm. Consider $q_{2}=0$ and let $q=q_{1}$. In this case, when $kp\ll 1\text{, and }q\ll\frac{1}{\sqrt{k\log\left(\frac{1}{kp}\right)}},$ the entropy is given by $H(X)=m\cdot k\cdot p\cdot\log\left(\frac{1}{kp}\right)+m\cdot k^{2}\cdot p\cdot q\cdot\left(\log k+\log\log\left(\frac{1}{kp}\right)\right)+1$. The authors show that their algorithm is order-optimal when, in addition to the above two conditions $kp\ll 1$ and $q\ll\frac{1}{\sqrt{k\log(\frac{1}{kp})}}$, we have $kp\ll m^{-\beta}\text{ for some fixed $\beta\in(0,1)$, and }1\ll kq$.
In Appendix D, we show that with only assuming $kp\ll m^{-\beta}$, the conditions of Corollary 4.11 hold and, the infected set is recovered with $O(H(X)+\mu)$ tests in expectation. In [ACÖ23], the authors prove $\mu\leq m\cdot k^{2}\cdot p\cdot(1/k+q)$. It is not hard to show that when $1/k\leq q$, which is a relaxed condition compared to $1\ll kq$, we have $\mu\leq H(X)$ and hence the dominating term in $O(H(X)+\mu)$ is $H(X)$, i.e. $O(H(X)+\mu)=O(H(X))$ and Corollary 4.11 thus provides an order-optimal performance guarantee. More generally, Theorem 4.10 provides the following result: Algorithm 1 can recover the infected set with probability $1-\epsilon$ using at most $O(H(X)+\mu/\epsilon)$ tests in expectation. This allows one to relax two additional conditions that [ACÖ23] has imposed, i.e. $kp\ll m^{-\beta}\text{ for some fixed $\beta\in(0,1)$, and }1\ll kq$.
6 Lower Bound and Optimality of Algorithm 1
In the probabilistic group testing literature, entropy $H(X)$ constitutes an order-wise tight lower bound [LCH+14] (see also Theorem 2.2.) The upper bounds we established in Theorems 4.8-4.10 and Corollary 4.11 involve not only $H(X)$ but also $\mu$, which is the expected number of infections. To shed light on the extent to which our proposed algorithm is order-optimal and our analytical upper bounds are tight, we seek to answer the following questions. (i) Is entropy always a tight lower bound in the generalized group testing problem with general correlation? (ii) Does $\mu$ play a fundamental role (lower-bounding the number of tests) or should one hope to tighten it in our results? (iii) Is the upper bound established in Theorem 4.10 and/or its corresponding algorithm order-wise optimal?
In this section, we present a family of graphs in which entropy is not a tight lower bound in Section 6.1, provide examples for when $\mu$ serves as a lower bound and when it does not in Sections 6.2 and 6.3, and derive conditions under which Algorithm 4.10 is order-optimal for random hyper-graphs in Section 6.4.
6.1 When Entropy is a Loose Lower Bound
By Theorem 2.2, we know that $H(X)$ is a lower bound on the number of tests. Consider the following example from [GLS22, Theorem 4] with some slight modifications to make the setting probabilistic.
Example 6.1.
Graph $G=(V,E)$ has $n$ nodes where each edge of size $n-1$ is included in $E$. To each each edge $e$, assign a uniform probability mass $p(e)=\frac{1}{n}$. Figure 2 demonstrates such a graph with $n=4$.
In the above example, the entropy is $\log n$. What makes this example important for us is that any algorithm with a constant probability of success requires at least $\Omega(n)$ tests. This can be proved using a very similar argument as in [GLS22, Theorem 4], and for the sake of completeness, we provide a short proof in Appendix B. What is significant here is that the gap between the entropy and the required number of tests is $O(n/\log n)$, which is much larger than expected. We can generalize the above example as follows:
Example 6.2.
Graph $G=(V,E)$ has $n$ nodes where at least $c$ fraction of all edges of size $d$ is included in $E$, for a constant $c\in(0,1)$. a uniform probability mass $p(e)=\frac{1}{|E|}$ is assigned to each edge $e$.
In Appendix C, we prove that any algorithm performing on Example 6.2 and achieving a constant probability of success needs to perform $\Omega(d)$ tests. These examples raise the question of finding other lower bounds or identifying classes of graphs where entropy is not a tight lower bound. Before proceeding further, suppose we run Algorithm 1 on this example. It is worth noting that for any constant $c$, the algorithm bypasses the initial stage and proceeds to test nodes individually in line 1, requiring $n$ tests. This example highlights scenarios where conducting individual tests is optimal, leaving no room for improving the theoretical guarantees in Theorem 4.10 and the following corollaries.
Now, we show that the Example 6.1 is not the only graph where the entropy is loose, and we generalize it for a larger family of graphs where $H(X)$ is a loose lower bound and at least $\mu$ tests are needed. Consider $n=|V|$, and let $m$, $d=|e|$ be given parameters. We build a graph with $n$ nodes and $|E|=m+O(d)$ edges where each edge has size $d$. Let the distribution $\mathcal{D}$ be uniform over the edges. Here, $H(X)=\log(m+O(d))$ and $\mu=d$. The graph is constructed in the following manner, where we start with an empty set for the edges and gradually add edges:
Example 6.3.
Graph $G=(V,E)$ and the distribution $\mathcal{D}$ are constructed as following:
  1. 1.
    Initiate $E=\emptyset$.
  2. 2.
    While ($|E|<m$): find a set $S\subseteq V$ with $|S|=d+1$ such that there is a subset of $S^{\prime}\subset S$ of size $|S^{\prime}|=d$ that is not currently in $E$. Add all subsets of $S^{\prime}\subset S$ of size $|S^{\prime}|=d$ to $E$, i.e.
    $E=E\bigcup\limits_{S^{\prime}\subset S,|S^{\prime}|=d}\{S^{\prime}\}.$
  3. 3.
    Set $\mathcal{D}$ to be the uniform distribution over $E$.
The first time that $|E|$ exceeds $m$, there are at most $d$ edges added, hence $|E|=m+O(d)$. We claim that at least $\Omega(d)$ tests are needed to recover the infected set. Suppose $e^{*}$ is the sampled edge, so at some point in line 2 $e^{*}$ was added to $E$. Hence there is a set $S$ of size $d+1$ that contains $e^{*}$ and there are $d$ edges inside of $S$. Now even if the algorithm knows the set $S$, using a similar argument as in [GLS22] and Appendix B, the algorithm still needs $\Omega(d)$ tests as only one node is negative and it is equally likely to be any of them.
Now if $\mu=d>\log(m+O(d))=H(X)$, then the entropy lower bound is loose and $\mu$ is an order-wise tight lower bound. This happens when $2^{d}>m+O(d)$.
6.2 When $\mu$ is not a Lower Bound
In all of the examples provided so far, either $\mu$ or $H(X)$ is a tight lower bound on the number of tests. One might conjecture that $\max(\mu,H(X))$ might be a lower bound on the number of tests. The following example shows that $\mu$ is not necessarily a lower bound.
Example 6.4.
Consider a graph $G=(V,E)$ comprising $n$ nodes and the edges are of the form $e_{i}=\{v_{1},\ldots,v_{i}\}$ and $p_{e_{i}}=1/n$ for each $i$.
Figure 5: An instance of the graph from Example 6.4 with $n=9$ nodes.
Here, we have $\mu=\frac{n+1}{2}=\Omega(n)$. But by a binary search algorithm, we can find the target edge in $O(\log n)$. More precisely, we first test $v_{n/2}$. If the result is positive, it means that the target edge is among $e_{i},i\geq n/2$ and the next step is to test $v_{3n/4}$. If the result is negative, then the target edge is among $e_{i},i<n/2$, and the next step is to test node $i/4$. By continuing this, we can find the target edge in $O(\log n)$ tests.
Note that for the above example, $H(X)$ is a tight lower bound. The following example provides a graph where not only $\mu$ is not a lower bound, $H(X)$ is a loose lower-bound too.
Example 6.5.
Consider a graph with vertex set $V=V_{1}\cup V_{2}\cdots\cup V_{n}$ where $|V_{i}|=n$, so $|V|=n^{2}$ and we call each $V_{i}$ a community. Community $V_{i}$ is composed of nodes $V_{i,j}$ for $j=1,\ldots,n$. The edge set is composed of “small edges” and “large edges”. The set of small edges is defined by $E_{1}=\{e\mid\forall j,k:e=V_{j}\setminus\{V_{j,k}\}\}$. So there are $n^{2}$ small edges and each small edge has size $n-1$. The set of large edges is defined by $E_{2}=\{e\mid\forall i:e=V\setminus V_{i}\}$, so there are $n$ of them and each has size $n(n-1)$. Here, $E=E_{1}\cup E_{2}$. The distribution $\mathcal{D}$ puts half of the mass on $E_{1}$ and half of the mass on $E_{2}$, and the edges in $E_{1}$ (respectively $E_{2}$) have the same probability among themselves. Figure 6 shows and example with $n=4$.
Figure 6: The graph in Example 6.5 with $n=4$, 16 nodes, and 20 edges. “Small” edges have black border, and “large” edges have red border.
In the above example, $\mu=\Omega(n^{2})$ and $H(X)=O(\log n)$. Following Example 6.1, it is not hard to see that $\Omega(n)$ tests are needed, because even after knowing whether the sampled edge is in $E_{1}$ or $E_{2}$, the algorithm needs to find the single negative node (in some community $V_{i}$) to recover the sampled edge.
But $O(n)$ tests are sufficient: one can pick a single node from each $V_{1},V_{2},V_{3}$ and test them individually. If at least two of them are negative, it means the sampled edge is in $E_{1}$, otherwise, it is in $E_{2}$. If the sampled edge is in $E_{1}$, using $\log n$ tests the infected $V_{i}$ can be detected and by testing all of its nodes, the sampled edge is recovered. If the sampled edge is in $E_{2}$, we can find the negative community using $n$ tests by testing each community as a whole. In either case, $O(n)$ tests is sufficient.
The above examples show that the lower bound does not necessarily obtained by simple properties of the graph, such as $\mu$ and entropy. Finding a tight lower bound is an interesting future direction.
6.3 When $\mu$ is a Lower Bound for Achieving Zero Probability of Error
We have already established that $H(X)$ is a lower bound. For any graph, if we show that $\mu$ is also a lower bound, then Algorithm 1 is order-wise optimal. One obvious example of such graphs is when $\mu\leq H(X)$, where Algorithm 1 would be order-wise optimal.
In this section, we provide a family of graphs where $\mu$ is not necessarily smaller than the entropy, but we prove that $\mu$ is a lower bound when the target probability of error is 0. The example is as follows:
Example 6.6.
Graph $G$ has this property that for every node $v$ and every realization of $V\setminus v$, there is a positive chance that $v$ is negative and there is a positive chance that $v$ is negative. In other words, for every $e\in E$ and $u\notin e$, there is an edge $e^{\prime}=\{v\}\cup e$ where $e^{\prime}\in E$.
Claim 6.7.
We claim that any testing algorithm on the graph in Example 6.6 with zero probability of error needs at least $\mu$ tests.
Proof.
First, we prove that every positive node $v$ should be presented in a test in order to achieve a zero probability error. If not, even knowing the states of $V\setminus v$, there is a positive chance that $v$ is positive or negative, which contradicts the assumption of zero error. Second, we prove that each positive node should be at least in a test where every other node in the test is negative. Suppose not, and for every test for a node $v$ that is positive there is another positive node. Then by changing the state of $v$ from positive to negative, the results of the tests remain the same. However having $v$ as a negative node corresponds to a valid edge by the assumption, contradicting the zero error assumption. This shows that for any graph with the realized edge $e^{*}$, at least $|e^{*}|$ tests are needed, meaning that on average, at least $\mu$ tests are needed. ∎
6.4 Random Graphs and Regimes of Optimality
Consider $d$-regular random graphs, where each subset $S\subseteq V$ of size $|S|=d$ is included in the graph with probability $r$, independent of the rest, to make the graph $G_{r}=(V,E_{r})$. Then $G_{r}$ is revealed with a distribution $\mathcal{D}$ defined on the edges. We analyze two regimes of interest by examining the upper and lower bounds in each case. The first regime is when $r$ is small and the graph is sparse, and the other is when $r$ is large and the graph is dense. The following shows that when $r$ is large, there is a distribution $\mathcal{D}$ over the edges $E_{r}$ so that $\mu$ is a lower bound.
Proposition 6.8.
Suppose graph $G_{r}=(V,E_{r})$ is given with the parameter $r\gg(\frac{d}{n})^{1/d}$. Then there is a distribution $\mathcal{D}$ on $E_{r}$ so that any algorithm with a constant probability of success needs $O(\mu)=O(d)$ number of tests on average.
Proof.
Suppose that $r\gg(\frac{d}{n})^{1/d}$. We partition the nodes into groups of size $d+1$, so there are $\frac{n}{d+1}$ groups. The probability that all the edges in one group are sampled is $r^{d+1}$, so with probability $(1-r^{d+1})^{\frac{n}{d+1}}$, no such group has all its edges sampled. Because $(1-r^{d+1})^{\frac{n}{d+1}}\simeq e^{-r^{d+1}{\frac{n}{d+1}}}$, if $r\gg(\frac{d}{n})^{1/d}$, this probability goes to zero. Hence, with high probability, there are $d+1$ nodes where every subset of size $d$ amongst them is a sampled edge. Call such a subset $S$ and all the the edges inside $e_{1},e_{2},\ldots,e_{d+1}$, where $\forall i:e_{i}\in E_{r}$ and $e_{i}\subset S$ and $|S|=d+1$. Now define the distribution $\mathcal{D}$ such that it only puts weights on $e_{1},e_{2},\ldots,e_{d+1}$, uniformly at random, i.e. $\forall i:\mathcal{D}(e_{i})=\frac{1}{d+1}$. Following the line of argument in Example 6.1, we conclude that any algorithm with a constant probability of success needs at least $O(\mu)=O(d)$ number of tests. ∎
Note that Algorithm 1 performs $O(H(X)+\mu)=O(H(X)+d)$ in the above regime by Theorem 4.8, and we already know that $H(X)$ is a lower bound. The above proposition shows that $O(d)$ is also necessary, proving that Algorithm 1 is order-wise optimal in this regime.
Proposition 6.9.
Suppose graph $G_{r}=(E_{r},V)$ is given with the parameter $r\ll(\frac{d}{n})^{O(d)}$. Then Algorithm 1 finds the infected set using $O(H(X))$ tests on average.
Proof.
We prove that when $r\ll(\frac{d}{n})^{O(d)}$, with high probability there is no subset of nodes of size at most a constant of $d$, let’s say $2d$ where more than a constant number of edges amongs them are sampled. If this is true, then when the Algorithm 1 gets to the Line 1 (recall that this is the line that the algorithm performs individual testing), there is only a constant number of edges remaining and it can test each of the edges instead of each of the nodes. Note that testing one edge here is possible using only one test. This is because the graph is $d-$regular and by testing $V\setminus e$, we can check if $e$ is the target edge or not.
The probability that more than $c^{\prime}$ edges, for a constant $c^{\prime}$, are sampled among $2d$ nodes is less than $r^{c^{\prime}}{2d\choose c^{\prime}}<(rd/(c^{\prime}/2))^{c^{\prime}}$. There are (${n\choose 2d}$ candidates for picking up a subset of size $2d$, and (${n\choose 2d}<(\frac{n}{2d})^{2d}$. Hence by union bound, the probability that there is a subset of size $2d$ with more than $c^{\prime}$ edges is at most $(\frac{n}{2d})^{2d}(rd/(c^{\prime}/2))^{c^{\prime}}$. When $r\ll(\frac{d}{n})^{O(d)}$, this goes to zero and hence with high probability, fewer than $c$ edges exist when Algorithm 1 gets to Line 1 and performs at most $c^{\prime}$ tests.
So with high probability, the total number of tests is $O(H(X))$ in the first phase of the algorithm on average, and $c^{\prime}$ tests in the second phase. So the total number of tests is $O(H(X)+c^{\prime})=O(H(X))$ on average, which shows that Algorithm 1 is order-wise optimal.
The above proposition shows that the analysis provided in Theorem 4.10 is worst-case scenario and even though we have a term $\mu=d$ in the upper bound of the theorem, Algorithm 1 is order-wise optimal.
7 Towards Non-adaptive Group Testing
Section 3 concerned adaptive group testing. In this section, we study the SNAGT model which is a model we proposed in Section 2 as a model in between adaptive and non-adaptive testing. To recall, in SNAGT, tests are designed beforehand and do not depend on prior test results. They are done sequentially and we know the result of a test just after performing it. While the test design does not depend on the result of the previous tests, we assume that we can stop testing at any point. In the following, we first focus on the SNAGT model, propose a testing strategy and analyze its performance. Then we discuss challenges one faces in designing fully non-adaptive algorithms.
As we already saw, when the testing is done adaptively, we get the best results when the number of infections is concentrated around its mean $\mu$. In other words, with high probability $|e^{*}|=\mu+o(\mu)$. In this section, we consider the same assumption, i.e., the size of the infected set is concentrated around its mean.
The following theorem shows that using $O(\mu H(X)+\mu\log n)$ tests, in the SNAGT mode, we can find the target edge with high probability.
Theorem 7.1.
Suppose for a graph $G$, the probability distribution $\mathcal{D}$ is such that $\Pr_{e\sim\mathcal{D}}(|\left|e\right|-\mu|>\epsilon\mu)=o(1)$ for any constant $\epsilon>0$. There is an algorithm that performs $O(\mu\cdot H(X)+\mu\log n)$ tests in expectation and recovers the infected nodes with an error probability that goes to 00 as $n\rightarrow\infty$.
Proof.
To simplify the proof here, we assume that every edge has the same size $d$. A very similar and more general proof is given in Appendix E.
We follow a proof technique similar to [LCH+14, Theorem 2]. In particular, we partition the edges into subgraphs so that the edge probabilities in the same subgraph are close to each other. Suppose the edges are partitioned into subgraphs $G_{1},G_{2},\ldots,G_{k}$ where $G_{i}=(V,E_{i})$ and $\forall e\in E_{i}:\frac{1}{2^{i}}\leq p(e)\leq\frac{1}{2^{i-1}}$. We can assume $k\leq n\log n$ if we want to have an error that goes to zero, as any edge in $e_{i},i>n\log n$ has probability at most $\frac{1}{2^{n\log n-1}}$ and there are at most $2^{n}$ of them, so the total error probability from this is at most $\frac{2^{n}}{2^{n\log n-1}}\rightarrow 0$. Denote $m_{i}=|E_{i}|$ as the number of edges in $G_{i}$, and $\mathbb{P}(G_{i})=\sum_{e\in E_{i}}p(e)$ the probability that the target edge is sampled from $G_{i}$.
First, imagine that we have the information that the target edge is in the $i$’th hypergraph. We perform the tests randomly, i.e. we pick each node with probability $1/d$ independently. Then, using [GLS22, Theorem 4], we can find the target edge with high probability using $d\log m_{i}$ tests. So on average, the number of tests is:
$\mathbb{E}[T]$ $=\sum_{i}\mathbb{P}(G_{i})d\log(m_{i})$
$=d\sum_{i}(\sum_{e\in G_{i}}p(e))\log(m_{i}).$
But $m_{i}\leq 2^{i}$ and $1/2^{i}\leq p(e)\leq 1/2^{i-1}$, so $m_{i}\leq 2/p(e)$ for all $e\in E_{i}$. Hence we get
$\mathbb{E}[T]$ $=d\sum_{i}(\sum_{e\in G_{i}}p(e))\log(m_{i})$
$\leq d\sum_{i}\sum_{e\in G_{i}}p(e)\log(2/p(e))$
$=d\sum_{e\in E}p(e)(1+\log(1/p(e)))$
$=d+dH(X).$
So in this case, we get $O(dH(X))$ tests in expectation.
But of course, we do not know which $G_{i}$ contains the target edge. To fix this, we propose the following algorithm: keep testing randomly as before, and rule out inconsistent edges in each subgraph until at least one subgraph has exactly one edge. Let $SG$ be the set of subgraphs with exactly one edge at this point. Then perform $2d\log n$ additional tests one by one and rule out inconsistent edges. If the single edge remains in exactly one subgraph after the tests, return the edge as the target edge. If there is more than one subgraph in $SG$ with exactly one edge, return one of them at random. If at one point during the $2d\log n$ tests, single edges are ruled out in all subgraphs of $SG$, continue the random test until another component has exactly one edge again and repeat. We proved that conditioned on the target edge being in the $i$’th subgraph, the algorithm performs $O(d\log m_{i}+d\log n)$ tests. Hence by the above argument, we perform $O(dH(X)+d\log n)$ tests in total.
Now we prove that the error probability goes to zero. Suppose the target edge is in $i$’th subgraph. We know that if the algorithm performs $O(d\log m_{i})$ tests, there is only one edge remaining in it with high probability and after the $2d\log n$ “sanity check” tests, it remains in the graph. Let $err_{1}$ be the event that the algorithm outputs an edge before performing $d\log m_{i}$ tests. Then it means the algorithm found one or more subgraphs with a single edge in them, and after performing $d\log n$ tests, these subgraphs are not ruled out. The probability that a non-target edge survives after $2d\log n$ tests is $1/n^{2}$ and there are at most $n$ of them, as the number of subgraphs is bounded by $n$ and the edge is the only one remaining in the subgraph. Hence by union bound, this probability is at most $\mathbb{P}(err_{1})<1/n$. Also, define $err_{2}$ be the event that after $O(d\log m_{i})$ tests, more than one edge remains in $G_{i}$. In [GLS22], they showed that this has a probability that goes to zero. Hence, the total probability of error goes to zero.
We have not been able to generalize the above technique to get general upper bounds for non-adaptive group testing. To recall, in non-adaptive testing, all the tests are performed at once. In this case, each test is independent of the previous ones, and only after seeing the results of all tests can we recover the infected set. Non-adaptive group testing has been extensively studied in the literature for independent group testing [LCH+14, WGG23, PR08], and typically, state of the art non-adaptive methods are differing from the adaptive counterpart only by $O(\log n)$ or $O(d)$ (multiplicative factor). For instance, $O(d^{2}\log n)$[DHH00] is needed in the classic non-adaptive combinatorial setting (compared to $O(d\log n)$ tests for adaptive) or $O(\log n(H(X)+\mu))$[LCH+14] in non-adaptive independent group testing (compared to $O(H(X)+\mu)$ tests for adaptive). Example 6.4 below shows that this is not necessarily true for general correlation and that the gap between adaptive and non-adaptive methods can be large, highlighting that the problem may be fundamentally more challenging than its counterpart in independent group testing settings.
Claim 7.2.
Any non-adaptive algorithm with probability error at most $\frac{k}{n},k\in\{0\}\cup\mathbb{N}$, needs to do at least $n-k-1$ tests in Example 6.4.
Proof.
Suppose we have the additional information that either $e_{i-1}$ or $e_{i}$ is the target edge, $2\leq n$. Between these two cases, only node $i$ has a different state. In order to correctly decide which one is the target edge, there must be a test where $i$ is the smallest index of the node that the test contains, otherwise, no algorithm can do better than random guessing. There is $2/n$ probability that $e_{i}$ or $e_{i-1}$ is sampled, and if no test contains $i$ as its smallest index, with probability at most $1/2$ it can recover the correct edge, hence incurring an error of $1/n$. There are $n-1$ candidates for the smallest index, and by not having one of them there would be an additional $1/n$ error. hence by missing $k$ of them, we will have a probability $k/n$ of error. ∎
Note that there is an adaptive algorithm for the above example with $\log n$ tests. This shows that the gap between adaptive and non-adaptive algorithms can be as large as $n/\log n$.
8 Noisy Group Testing
Up until now, we have assumed that the tests are noiseless, i.e. the test is positive iff there is at least one infection in the group that is tested. In this section, we extend our results to the case of noisy tests. We consider the symmetric noisy model introduced in Section 2.1, where each test result is flipped with a small probability $20<\delta<1/$ 2, independent of the other tests.
A simple and standard argument shows that if we repeat each test for $T_{R}>\frac{\log(\mu+H(X))}{1-4\delta(1-\delta)}$ times, and then take the majority, with high probability, we get the correct result for all the tests. So we incur an additional factor of $\frac{\log(\mu+H(X))}{1-4\delta(1-\delta)}$.
Another way to solve this issue is to calculate the posterior probability considering the fact that tests are noisy. The following theorem shows that with small modifications in Algorithm 1, we can find the infected set better than just repeating each test.
Theorem 8.1.
Suppose the result of each test is flipped with probability $\delta<1/2$, $\delta<\epsilon$, and $c>1-\frac{1}{(1-\delta)^{1-\delta}\delta^{\delta}}$. There is an algorithm that performs $\mathbb{E}[L_{1}+L_{2}]\leq\frac{1}{\log(1-c)-H(\delta)}H(X)+\frac{\log\mu}{1-4\delta(1-\delta)}\frac{\mu}{2\epsilon(1-2c)}$ tests in expectation and returns the correct edge $e^{*}$ with probability $1-2\epsilon$.
Proof.
We first describe how we update the posterior when the tests are noisy. Consider an edge $e$ and a test $T\subseteq V$, where $e\cap T=\emptyset$. If $T$ is negative, there are two possibilities. Either it is not flipped, which happens with probability $1-\delta$ and $p(e)$ is updated to $\frac{p(e)}{\sum_{f\in E^{\prime}}p_{f}}$, where $E^{\prime}$ is the set of edges that have no intersection with $T$, or the test is flipped and $p(e)$ is updated to 00. So the posterior probability for $e$ is $p(e)=(1-\delta)\frac{p(e)}{\sum_{f\in E^{\prime}}p_{f}}$ Similarly, if the test is positive, if it is not flipped $p(e)$ is updated to 00 and if it is flipped, $p(e)$ is updated to $\frac{p(e)}{\sum_{f\in E^{\prime}}p_{f}}$, hence the posterior probability is $\delta\frac{p(e)}{\sum_{f\in E^{\prime}}p_{f}}$. With the same argument, we can obtain the posterior probability when $e\cap T\neq\emptyset$, and only the posterior for positive and negative tests are swapped.
Again by Markov Bound, with probability $1-\epsilon$ the number of infections is less than $\mu/\epsilon$. From now on, we allow for $\epsilon$ error, as $(1-\epsilon)^{2}>1-2\epsilon$, which is our target error. Now, instead of line 1 where the algorithm returns the single edge, if there is an edge with probability more than $1-\epsilon$, the algorithm returns it. Suppose $e^{*}$ is the target edge and $T$ tests are performed in the first phase. Then with high probability and by the new posterior obtained in the last part, we have $q_{e^{*}}=(1-\delta)^{(1-\delta)T}\delta^{\delta T}(\frac{1}{1-c})^{T}=((1-\delta)^{1-\delta}\delta^{\delta}\frac{1}{1-c})^{T}>(c^{\prime})^{T}$ where $c^{\prime}>1$. Hence it takes $\log{1/p(e)}/\log c^{\prime}$ tests for $q_{e^{*}}$ to become greater than $1-\epsilon$. Similar to Corollary 4.6, the number of tests in this phase is $\frac{1}{\log c^{\prime}}H(X)$
Now for the second phase, same as the previous part, we repeat each single test in line 1 and take the majority. Again, if we repeat each test $O(\frac{\log\mu}{1-4\delta(1-\delta)})$ times, with high probability we recover all the nodes correctly and hence we perform $\frac{\log\mu}{1-4\delta(1-\delta)}\frac{\mu}{2\epsilon(1-2c)}$ tests. ∎
Theorem 8.1 shows that our framework can handle the noise and update the posterior accordingly. The following theorem shows under the same definition of noise, we can achieve a result in the SNAGT setting by just repeating the tests.
Theorem 8.2.
Suppose the result of each test is flipped with probability $\delta<1/2$, and for a graph $G$, the probability distribution $\mathcal{D}$ is such that $\Pr_{e\sim\mathcal{D}}(|\left|e\right|-\mu|>\epsilon\mu)=o(1)$ for any constant $\epsilon>0$. Then, there is an algorithm in the SNAGT setting that performs $O(\frac{\log[\mu\cdot H(X)+\mu\log n)]}{1-4\delta(1-\delta)}[\mu\cdot H(X)+\mu\log n)]$ tests in expectation and recovers the infected nodes with an error probability that goes to 00 as $n\rightarrow\infty$.
Proof.
We employ the same algorithm described in Theorem 7.1, but modify it to address noise by incorporating test repetition and majority voting. Specifically, each test is repeated $\frac{\log[\mu\cdot H(X)+\mu\log(n)]}{1-4\delta(1-\delta)}$ times, and the outcome of the test is determined by the majority vote among these repetitions.
By choosing the number of repetitions as above, we ensure that the probability of a single test producing an incorrect majority outcome is at most $\mathrm{err}\ll\frac{1}{\mu\cdot H(X)+\mu\log(n)}.$ Applying a union bound over all tests, the probability that any test is incorrect tends to zero as $n$ grows.
Consequently, with high probability, we may assume that all tests are effectively noiseless. Under this assumption, the situation reduces exactly to the noise-free setting of Theorem 7.1. Hence, the same performance guarantees and conclusions from that theorem apply, completing the proof. ∎
9 Conclusion and Future Work
In this paper, we developed a novel framework for group testing with general statistical correlation among nodes. Our model considers a probability distribution on all possible candidate subsets of nodes, captured by a hypergraph. We developed group testing algorithms that exploit correlation in two settings, namely adaptive and semi-non-adaptive group testing. In the adaptive setting, our algorithm consists of two stages: informative testing and individual testing. We demonstrate that the number of tests required is a function of the entropy and the expected number of infections, achieving order-wise optimality when the expected number of infections is less than the entropy. In particular, when the number of infections is concentrated around its mean, the number of required tests is upper bounded by $O(H(X)+\mu)$, in expectation. We provided several examples illustrating both the worst-case optimality and limitations of the algorithm. The algorithm remains robust in noisy settings (with symmetric noise), as it can correctly update its belief despite errors.
We further introduced and studied a semi-non-adaptive setting– a model in between adaptive and non-adaptive settings. In this model. tests are designed independently from prior test results (similar to non-adaptive group testing), but it is assumed that testing can stop once the infected set is found (sequentially as a function of prior test results). In this setting, we provided a group testing strategy that requires at most $O(\mu H(X)+\mu\log n)$, in expectation. The algorithm can be extended to the noisy group testing setting with symmetric noise.
While we discussed conditions under which our algorithm and its corresponding performance upper bound are near-optimal, it remains open whether our algorithm has optimal performance in general. Future work will focus on providing tighter lower and upper bounds for both adaptive and semi-non-adaptive settings, exploring the role of the correlation structure. Additionally, we aim to explore the design of non-adaptive algorithms despite the challenges highlighted in this paper. This includes designing such algorithms for general graphs or specific families of graphs that could significantly reduce the number of tests needed. Furthermore, we will continue to investigate new correlation models that could expand the practical applicability of our framework, ultimately enhancing our ability to perform efficient group testing in diverse, real-world environments.
Appendix A Modeling Statistical Models with Hypergraphs
Here, we give explicit formulas to compute the equivalent hypergraphs that model previous works.
As discussed in Section 2.2, In [NKC+22] they consider a simple graph $H=(V_{H},E_{H})$, and random graph $H_{r}$ is obtained by keeping each edge with probability $r$. Then, every component is infected with probability $p$, i.e. all the nodes inside are infected, or not infected with probability $1-p$, i.e. all the nodes inside are not infected. In order to compute $\mathcal{D}(S)$, the probability that exactly set $S\subseteq V_{H}$ is infected, in graph $H_{r}$, set $S$ should be disconnected from the rest. Let $\mathcal{H}_{S}$ be such graphs. Let $C(H_{S})$ be the number of components on graph $H$ induced by $S$, and $E(H_{S})$ be the number of edges in $H_{S}$ induced by $S$. The $\mathcal{D}(S)$ is computed as follows:
$\mathcal{D}(S)=\sum_{H\in\mathcal{H}_{S}}p^{C(H_{S})}\cdot(1-p)^{C(H_{V\setminus S})}\cdot r^{E(H_{S})+E(H_{V\setminus S})}\cdot(1-r)^{E(H_{V})-E(H_{S})-E(H_{V\setminus S})}.$
In [AU23], they have the same model except only one component is infected uniformly at random. Define $\mathcal{H_{S}}$ be the graphs where $S$ is a connected component, disconnected from the rest of the graphs. Then $\mathcal{D}(S)$ is computed as:
$\mathcal{D}(S)=\sum_{H\in\mathcal{H}_{S}}\frac{1}{C(H_{V})}\cdot r^{E(H_{S})+E(H_{V\setminus S})}\cdot(1-r)^{E(H_{V})-E(H_{S})-E(H_{V\setminus S})}.$
For [ACO21], we can compute $\mathcal{D}(S)$ by taking the sum over all possible seeds. Let $N(v)$ be the number of seeds with the same community as $v$ and $n$ be the number of nodes. Then:
$\mathcal{D}(S)=\sum_{T\subseteq S}q^{T}\cdot(1-q)^{n-T}\cdot\prod_{v\in T\setminus S}[1-(1-q_{1})^{N(v)}(1-q_{2})^{|T|-N(v)}].$
Appendix B $\Omega(n)$ Tests are Needed in Example 6.1
Theorem B.1.
Any algorithm with a constant probability of success with the graph in Example 6.1 requires at least $\Omega(n)$ tests.
Proof.
First, note that every test of size two or greater will yield positive results, rendering them redundant. Consequently, the problem is effectively simplified to individual testing when a single individual is found to be negative.
Every test is positive unless we find the negative node, so it takes $\Omega(n)$ tests to identify that particular node if we want to have a constant probability of success.
Appendix C A Proof for Example 6.2
Consider all the subsets of $S$ where $|S|=d+1$. We prove that the average number $E(S)$ is $c(d+1)$. There are (${n\choose d+1}$ of subsets of size $d+1$, and each edge appears in $n-d$ of such subsets. Hence the average number of edges inside $S$ is $\frac{(n-d)|E|}{{n\choose d}}=\frac{c(n-d){n\choose d}}{{n\choose d+1}}=c(d+1)$.
Now if the algorithm has additional knowledge that the target set is in a subset $S$ where $|S|=d+1$ and $E(S)\geq c(d+1)$, it means there are $c(d+1)$ candidates for the single noninfected node, hence any algorithm needs to perform $O(d)$ tests for a constant probability of success.
Appendix D Proof of [ACÖ23] From Section 5
Let $C_{i}$ be a random variable that is $1$ when community $i$ is infected and $S=\sum_{i}C_{i}$. We show that under the condition $kp\ll 1$, $S$ is unlikely to get values much greater than its mean. We have $\mathbb{E}[S]=m\mathbb{E}[C_{1}]=m(1-p)^{k}\simeq me^{-pk}=e^{\log m-pk}$. We know that $C_{i}$’s are independent of each other, hence by Chernoff bound we have
$\mathbb{P}(S>(1+\delta)\mathbb{E}[S])<e^{\frac{\mathbb{E}[S]\delta^{2}}{2}}=e^{\frac{e^{\log m-pk}\delta^{2}}{2}}.$
Now because $pk\ll 1$, the above probability is $O(e^{m\delta^{2}})$. By having large constant $\delta$, we have $S=O(\mathbb{E}[S])$ with high probability.
Now that the number of infected communities is concentrated, we argue that the number of infections in a community is also concentrated. Note that when $kp\ll m^{-\beta}$, we only see at most a constant number of seeds in each community (with high probability). Since every non-seed node is exposed to a constant number of seeds, the probability that a non-seed node becomes infected is $O(q)$, and hence the expected number of infections in each community is $O(kq)$ and the expected number of non-seed infections is $O((n-m)q)$ and the non-seed nodes are independent of each other, i.e. one being infected or not does not affect other nodes. Now by Chernoff bound the probability that the number of infections is greater than $2(n-m)q=O(kmq)\gg m$ is $o(1)$. This shows that the total number of infections is not more than a constant factor greater than its mean, and hence the condition of Corollary 4.11 holds.
Appendix E Proof of Theorem 7.1
We already proved the theorem when every edge has size $d$. We are using the same algorithm as described for the $d-$regular case, where each node is in the test with probability $1/\mu$ at random, and slightly change the analysis. We first prove that when the target edge is in $i$’th subgraph, $\mu logm_{i}$ tests are enough to detect it. The argument is very similar to [GLS22]. Let $\ell=\mu-c$ a and $u=\mu+c$ where $c=o(\mu)$ and $\Pr_{e\sim\mathcal{D}}(\ell<|e|<u)=1-o(1)$. We only focus on the case when $\ell<|e|<u$. For any pair of edges $e_{1}$ and $e_{2}$ in $G_{i}$, we call a test $T$ separates them if $T$ contains at least a node in $e_{1}$ and no node in $e_{2}$, or $T$ contains at least a node in $e_{2}$ and no node in $e_{1}$. The probability that a random test separates $e_{1}$ and $e_{2}$ is at least $\frac{(1-1/\mu)^{u}}{\mu}=\Theta(\frac{1}{\mu})$. There are (${m_{i}\choose 2}=O(m_{i}^{2})$ different pairs, so if we do $3\mu\log m_{i}$ tests, by union bound, we can separate any two edges with high probability. This in turn means that we can rule out all the edges that are not the target edge. Hence, by knowing that the target edge is in $i$’th subgraph we can recover it using $\mu\log m_{i}$ tests. Using a similar argument, in expectation we perform $\mu H(X)$ tests in expectation when we know the subgraph that contains the target edge.
Now we perform random testing until there is at least one subgraph with exactly one edge and we perform $2\mu\log n$ additional test to verify it. Using a similar argument as above, the probability that the target edge remains after the additional tests is $O(1/n^{2})$ and there are at most $n$ subgraphs, hence the probability that we output it at any point is at most $O(1/n)$. This completes the proof